Current debt to equity ratio is .6 and its total capital is 800 bn. what is its debt to capital ratio ? and the dollar value of debt ?
Answer #1

Debt/Equity=0.6
=>(Debt+Equity)/Equity=(0.6+1)/1
=>Equity/(Debt+Equity)=1/1.6

Subtracting both sides by 1, we get;
=>1-Equity/(Debt+Equity)=1-1/1.6
=>Debt/(Debt+Equity)=1-1/1.6=.6/1.6
So, weight of debt=.6/1.6

Dollar value of debt=(.6/1.6)*(800)=300 bn

Debt to total capital ratio=300/800=0.375